OpenStudy (anonymous):

Two number are in ratio5:3 and their difference is 10. what is the product of two number? A. 250 B. 375 C. 175 D. 625

OpenStudy (roberts.spurs19):

Ok so if we call one number x and one number y we can reach a system of simultaneous equations as the difference between the two numbers is 10 we can say that x+10=y

OpenStudy (roberts.spurs19):

This means that we are assuming that y is the larger of the 2 numbers. This gives the other equation to be 3y=5x

OpenStudy (roberts.spurs19):

Do you know where to go from here?

OpenStudy (roberts.spurs19):

Hint: Solve the equations by substitution

OpenStudy (anonymous):

i have no idea which option is right in their options

OpenStudy (roberts.spurs19):

No worries! Ok so what we can do is say that 3(10 +x) =5x

OpenStudy (roberts.spurs19):

When we solve this we get that x=15 as y=10+x y must be 25

OpenStudy (roberts.spurs19):

When they say find the product, they mean multiply, therefore to get the answer you do 15*25 = 375

OpenStudy (roberts.spurs19):

Does this all make sense?

OpenStudy (anonymous):

thanks dud

OpenStudy (roberts.spurs19):

yw!